Fermi National Laboratory
 

Fermilab Shutdown Report for August 7, 2009 - August 14, 2009

 

Fermilab Shutdown Report

8/7/09 – 8/14/09

LINAC

Shutdown Week 9 – On Schedule

  • Preaccelerator
    • H - is operational putting out 56 mA
    • Continue slowly conditioning the I - system
    • Worked on trim package PS chassis
    • Supported PR and tour events
  • Low Energy
    • Replaced switch tubes in LRF2
    • Replaced 7651 tube in LRF1 driver
    • Worked on the LRF4 modulator
      • Station is now stable after an optical card was replaced. The spurious trips of the modulator ceased.
    • Worked on the QPS104 power supply
    • Tune up stations for better performance
  • High Energy
    • Completed work on the HE QPS
    • Initiated klystrons turn on
    • Ongoing work on cleaning the klystrons heat exchanger
      • Currently being serviced skid #1
      • Operational skids: #2 & 3
  • Next Week
    • Support NTF patient treatment
      • (Tuesday, Thursday, Friday)
    • - Address problems as they reveal
    • - Run 400MeV beam to the dump

Booster

Shutdown Week 9 – On Schedule

  • Tunnel Activities
    • Testing new corrector power supplies
    • GMPS work continues 95% complete
    • BRF ramping in progress
    • Pulsing kickers
    • Laser Profile Monitor testing
    • Re-core doors and S&S
    • FESS repaired LP-2 LCW pump at CUB
    • FESS replacing a section of the chilled water line Booster East Gallery 0600-1200
  • Next Week:
    • Today’s plan:
      • Safety system tests
      • GMPS hipot
    • Weekend plan:
      • GMPS commissioning
      • Power supply checkout
    • Next week’s plan
      • Power supply turn-on, repairs
      • Limited access time (controlled access)
      • Beam running on following weekend possibly earlier if all goes well

Antiproton Source

Accumulator 4-8GHz DP Kicker Upgrade: ~95% complete

The A50 bake ended this morning. Assuming no internal electrical problems, reconnection of exterior electronics and cooling water will be done on Monday. Power testing the tanks will follow.

Debuncher Septum Magnet Replacement: complete

Magnet is being pulsed during the dayshifts while we are here to monitor the vacuum. Today the shielding steel is being restacked.

Other Work

  • At this time, Pbar rings and all beamlines are closed up and under vacuum.
  • RF system work. Maintenance work on various RF systems continues. We will start testing some systems next week.
  • LCW leak repairs are almost complete.
  • Target station work continues. Inspection photos of the target will be taken next week.
  • LN2 flow to the A60 cryo pickups has been re-established.
  • AP2 vacuum was slow to recover. It was discovered that several ion pumps needed re-termination of their HV connections because of radiation damage to the cables and some sparking due to high humidity.
  • Vacuum at D:ISEP is poor. No external leaks.
  • We’ll be ready to start safety system tests on Monday 8/24 if possible.
  • When safety system testing is done, secure and start power supply testing.
  • We will stay in controlled access mode after 8/24.

Main Injector

Tunnel Work

MI-8 Multiwires

  • All Beam Valves relocated.
  • All Multiwires installed.
  • Texas Multiwire leaks fixed in place (no wires changed)
  • Baking 810, 811, 813 & 814 wires. May want to bake 830, 839 & 840

MI-40

  • Pipe fitters are done.
  • First kicker is in the tunnel, three of the six kickers are at MI-60 ready to start installation today. Two kickers are being tested at E4R. Seventh and final kicker is assembled and ready for testing.
  • Lead has been removed from the Lambertson. Two ion pumps replaced and leak is successful. Preparing for bake.
  • Installed 2 BPMs @ 403.
  • Ceiling plates have been installed for shielding.

MI eCloud (MI-52)

  • Installation complete.
  • Cabling in progress.
  • Baking absorbers.

Other Work

  • Installed shielding plates at MI-10.
  • Lead shielding has been removed from MI-10 and MI-40
  • All contractor tunnel work is complete.
  • MI-30 collimator LVDT investigation continues
  • MI-30 collimator mask installation scheduled to begin Aug. 24 th
  • MIPS & MIRF work going well, on schedule.
  • Installed HLS at MI-10.
  • LCW fill will begin after MI-40 Lambertson bake is finished. .

Next week

Finish bakes for MI-8 line and MI-52

Finish bake on MI-40 Lamberston

Install remaining kickers at MI-40 and complete vacuum work

Prepare for collimator mask work

Final Week

Tie up loose ends

Install collimator masks

Alignment

Recycler

Recycler

  • Should be able to close up Pelletron by August 19
    • New 2” GVM installed
      • ‘Hot spare’
    • Shaft bearing inspected
      • Cavities containing the Geo-rotors and belts have been cleaned
    • Transfer skid maintenance almost complete
      • Ordered oil for compressor
    • Install and test modified modulator today
  • Installed new ACCELNET computer in MI-31
    • ‘Hot spare’
  • Water/Glycol system cleanup and upgrade ( Complete)
  • NMR probes replacement ( Complete)
    • Shielding needs to be reinstalled
  • RR129 & RR400 (80% complete)
    • RR129 pot complete
    • Set back with new vials installation
      • Needed to drain flourinert at RR129 and wait for parts for RR400
  • Move 6 magnets ( complete)
  • Vacuum maintenance
    • Looks good at 401-409
    • Leak checked 100-103 & sublimated TSPs
      • No leak was found
    • New: Will bake IPs in the 800 line
      • Then sublimate TSPs
    • Start pump down of MI-10 region… Monday ?
      • ~Two weeks behind
  • Power supplies maintenance (90% complete)
  • VDPA1 & VDPA2: waiting for MI injector water to check for hose leaks
  • CIA crate activities ( Complete)
  • HLS installation at MI-10
    • All remaining work in service building
      • Except for troubleshooting
  • 1-2 GHz stochastic cooling system converted back to longitudinal
  • Latest complication/issue
    • Repositioned cable tray at MI-40 makes alignment surveys of Recycler elements difficult
      • 5 quads, 1 kicker, Lambertson and associated BPMs
      • Surveyors use side of the magnet fiducials (4) for which they don’t have a line of sight with the laser anymore
    • A solution is being worked out…

Tevatron

Other Items

  • Bad stands 10 of 128 done finish this Saturday
  • A1, A2 Hi Pot OK
  • Replaced D 11-3 re alinged D11-2 to 11-5
  • Still have beam tube vacuum in D1 to check
  • Spoiled vacuum in E-1
  • Vertical crystal collimator baking install next week
  • Still to do Dump resistor at F46
  • Finish HLS at B0
  • No alignment of low beta quads

NuMI

Shutdown Week 9

  • Install Target NT-03
    • Move Target NT-03 from MI8 to MI65
    • Rig Target Down Shaft into Target Hall
    • Move Target into Work Cell
    • Move Target Module into Work Cell
    • Repaired Z Drive Shaft Coupling
    • Reinstalled Z Drive into Module
    • Attach Target to Module
    • Complete Pressure Testing on Target
    • Survey Target NT-03 in Work Cell
    • Move Target & Module into Target Pile
    • Reinstall Blue Block and T-Block Shielding
  • Continue Installation of AH Dehumidification System
    • Complete Installation of Controls System
    • Complete Installation of Pumps & Heat Exchanger
    • Electrical Check Out Complete
  • EPB Dipoles PEEK Installation Continues
    • HV101-4, 101-5 & HV101-3 Survey Completed Today
  • New Texas Multi Wires for PM114 & PMTGT
    • Vacuum problems causing PM’s not to be installed
  • Install New Access Door to Emergency Walkway on Target Hall Side
  • Complete Low Power Anti-Neutrino Mode Testing on Horn Power Supply
  • Complete Hadron Monitor Installation
    • Emergency stairs reinstalled & electrical in-process
    • Started re-wiring of Hadron Monitor

Cryogenic System

TeV at LN2 Temperatures

Tevatron Operations and Maintenance Status

  • TEVATRON FRIG
  • Continued support AD/MSD for vacuum repairs: D3, E3, and Pbar Accumulator tanks are back at LN2 temperature. A1, A2, and D1 under LOTO for repairs by MSD. A3 is ready for cooldown but awaiting A2 repairs. F1 is cooling down to LN2 temperature. E1 is warming up to room temperature for repairs by MSD.
  • LN2 inventory control and controlling magnet temperatures at LN2.
  • Expanders group: Expander u-tubes are being re-stung back into service. Enormous backlog of shop repair and overhaul work.
  • Electrical group: Continuing upgrade of Cutler Hummer 120V power. Continuing maintenance of cryo I/O crate, electrical equipment cabinets, various low voltage electrical/electronic maintenance.
  • Compressor group: Mycom Compr relief valves replaced (6/8). 5/5 IHI cold comprs have been overhauled.
  • Ongoing numerous FRIG/Compressor controls/instrumentation maintenance/repairs.
  • Vacuum pumping of various cryo equipment in FRIG bldgs to start next week.
  • ---------------------------------------------------------------------------------------
  • Pending RT houses and relief maintenance, Tevatron cryo is on schedule with purification starting week of 8/24.

CHL HELIUM and NITROGEN PLANTS

  • CHL HELIUM and NITROGEN PLANTS
  • 99% of all scheduled routine mechanical and electrical maintenance is complete
  • Critical items:
  • Scheduled recycle TSI 2.5MW compressor overhaul in progress by outside subcontractor; test run today.
  • Expect delivery of repaired turboexpanders for the CHL coldbox-2 on Aug 17. Following by installation repaired turbines into coldbox-2, test run, and subsequent full cooldown with coldbox-2. Backup plan is to restart with coldbox-1 that operated before shutdown.
  • Scheduled compressor “B” and “D” cylinder valve maintenance completed by outside subcontractor.
  • Replacement of the plate&frame CHW/ICW heat exchanger done.
  • Repairs of the Nitrogen plant Marley water tower-subcontractor completed.
  • Setting up Mobile Helium Purifier at CHL for Tevatron FRIG purification.
  • -----------------------------------------------------------------------------------------------------
  • Pending successful TSI and coldbox-2 repairs, maintenance, CHL cryo is on schedule and planning resumption of 24/7 shiftwork week of 8/17.

Controls

Completed This Week:

  • TLG reboot
  • Startup support
  • General Controls Maintenance & Support

Next Week:

  • Possible loss of internet & email access from and VPN access to AD network Saturday (8/15)
  • Due to power outage in Wilson Hall
  • MAC Room Cooling
  • Support startup activities
  • General Controls Maintenance & Support

MiniBooNE

Everything is going well.

MTA

Magnet ready to be installed

FESS

Electrical and High Voltage

  • KAUTZ RD SS – buss duct & breaker PM’s WIP – 10yr PM work on the three SF6 345kv breakers progressing. One of them requires rebuild. Completion of work by 8/21.
  • MR & MI - air switch PM’s on schedule. Finding on FDR 23 compad – cracked insulator.
  • Switch yard – oil switch replacement start on Mon 8/17 – will have generator backup [FESS]

Mechanical

  • E0 – 75hp pond pump replaced.
  • MI 60 – glycol cooling line leak repair scheduled.
  • F-3 – by-pass pond pump replacement to be scheduled.
  • BGE – 2” chilled water line replacement [pipe chase] complete by Fri pm.

HVAC

  • MI construction areas – will return to check Marvair HVAC units & tunnel de-humidifiers.

ICW

  • Zebra Mussel treatment – completed on 8/8 & went very well – no problem areas.
  • Spot flushing – at Meson Line & strainer check at Switch yard set for Tues 8/25.
  • Valve 224 on Meson Line – replacement scheduled for Tue 8/18. [Flow reconfigurations]

CUB

  • 55LCW chiller[#3] - tube/header on schedule for replacement & on line by 8/31.
  • Process chiller #5 – requires plugging of 4 tubes – on line by end of next week.
  • DI regeneration column – plugged chemical distributor tubes replaced – backed up on regenerations into next week.

Other

  • Minos – glycol chilled water line repair scheduled for Thurs 8/20 [no cooling/dehum].
  • Wilson Hall shut down for electrical maintenance Sat 8/15. Critical systems on generators. Turbine Generators will require ~ 1 hr or longer shutdown in the afternoon has needed.

Switchyard 120 – Work has not started

Collider Detector at Fermilab – On schedule

D-Zero– On Schedule

Mechanical – On Schedule

Instrumentation– On Schedule

Alignment

Alignment

  • Survey of the berm for radiation shielding at MI10 after the civil construction is complete.
  • Survey of the berm for radiation shielding at MI40 after the civil construction is complete.
  • Align 7 new kicker magnets to be installed in the MI40 area. See Dave Capista or O'Sheg Oshinowo for details.
    • Working on today
  • As-found all the Main Injector and Recycler components in cells102-109. The essential components are: Recycler = magnets and cooling tanks no trims. 40 Components
  • As-found all the Main Injector and Recycler components after construction back fill in cells 341-404. Align all kickers and new beam pipe. 40 Components
  • Realign 6 EPB dipoles @ (HV101 magnet string).
    • Rough Align Completed
  • Align 3 replacement Fermi multiwires at PM101,PM105,
    • Rough aligned
  • PM114. 3 Components Not in place
  • Re-align Fermi multiwires at PM121 after stand adjustment (and re-check adjacent instrumentation).
    • New Texas Multiwires Not installed yet
  • Target Alignment on Monday
  • Stickmic survey of Muon system including CF and EF's after closing detector. The start and finish dates are preliminary and depend on shutdown progress.Shutdown 2009
  • Align 2 epb dipoles in enclosure B (VH94) before reworked epb dipoles. Align 2 new dipoles.
  • Align new permanent magnet to be installed on the Linac side. See O'Sheg Oshinowo or Carol Johnston for details.
  • Replace/ Install SEM AP0

Safety

Safety

  • 9.74 person-rem of absorbed dose first eight weeks
  • Shutdown ALARA plans are progressing very close to as planned
  • Highest single person dose 200 mrem
  • Individual dose rates are averaging less than 25 mrem per week
  • Projecting out to the end of the shutdown, total dose to individual workers
    should be under 300 mrem
  • Three radioactive dumpster violations
    • All three appear to be from contractors
  • Accelerator restart sheets are available from Mike Gerardi

Interlocks

  • Single Line Electrical Drawing Verifications
    • P-Bar 95% complete
    • Tevatron A Sector 20% complete
  • Audio warning system audio level control installation
    • 60% complete
  • Misc. Tevatron Interlock repairs & ground fault diagnostics installation
    • 70% complete
  • MINOS Absorber interlock modifications
    • 30% complete
  • Safety System Testing
    • VTS interlocks tested this past week
    • Booster enclosure testing completed this morning
    • MI-8 Line testing in progress
    • Booster CDC, Logic Module, and chipmunk testing later today

Shutdown Safety

  • Injuries
    • Two back strains in the past week. One First Aid case the other is a DART case.

Continue using the Take Five principles before, during, and after work activities to:

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/14/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory